OpenStudy (anonymous):

integrate (x)((x-2)power of one over 2)

OpenStudy (anonymous):

Is this what you intend?\[\int \frac{x}{\sqrt{x-2}}dx\]

OpenStudy (anonymous):

For starters, substitute \(u=x-2\) and don't forget to replace the \(dx\) as well as the \(x\) with expressions in terms of \(u\).

OpenStudy (anonymous):

no... x multiply by....

OpenStudy (anonymous):

I believe yak's idea will still work in that case.

OpenStudy (anonymous):

\(\displaystyle \int x \sqrt{x-2}\ dx\)?

OpenStudy (anonymous):

joemath314159 is right, the same substitution applies.

OpenStudy (anonymous):

can you show it???

OpenStudy (anonymous):

let:\[u=x-2\Longrightarrow x=u+2\]Then we also have:\[u=x-2\Longrightarrow du=dx\]so our integral becomes:\[\int\limits x\sqrt{x-2}dx=\int\limits (u+2)u^{\frac{1}{2}}du=\int\limits u^\frac{3}{2}+2u^{\frac{1}{2}}du\]

OpenStudy (anonymous):

tq bro...nice,